OpenStudy (acacia):

Find the determinant of [6 3] [10 5] umm what's a determinant?

jimthompson5910 (jim_thompson5910):

in the case of the 2x2 matrix  the determinant of that is D = a*d - b*c

jimthompson5910 (jim_thompson5910):

so in your case, you have [6 3] [10 5] which means that a = 6, b = 3, c = 10 and d = 5 so, D = a*d - b*c D = 6*5 - 3*10 D = 30 - 30 D = 0 So the determinant is 0

jimthompson5910 (jim_thompson5910):

the determinant is important for a lot of applications, but you'll find it's really useful in determining if you will have a consistent system or not

jimthompson5910 (jim_thompson5910):

yes pretty much, that's a good way to remember it

OpenStudy (acacia):

Ohh okay! Gee thanks smarty pants!!

jimthompson5910 (jim_thompson5910):

lol you're welcome

OpenStudy (acacia):

@jim_thompson5910 can you help me find the inverse of another problem?

jimthompson5910 (jim_thompson5910):

sure, what do you need help with?

OpenStudy (acacia):